# Robust Attitude Control of Nonlinear Multi-Rotor Dynamics with LFT Models and $\mathcal{H}_{\infty}$ Performance

### v1

@@ I Introduction
A key advantage of robust control frameworks lies in their compatibility with Linear Parameter-Varying (LPV) modeling, which enables exact representation of nonlinear dynamics without local linearization or approximation [12]. The rotational dynamics of multi-rotors contain trigonometric and rational expressions that can be modeled precisely using static nonlinear operators, which are then treated as structured, bounded uncertainty blocks within the Linear Fractional Transformation (LFT) formalism. This representation maintains the full fidelity of the nonlinear dynamics while enabling systematic controller synthesis via $\mathcal{H}_{\infty}$ technique.
@@ I-A New Contributions
This work advances the application of structured uncertainty frameworks in multi-rotor control by integrating LPV-LFT modeling with robust synthesis techniques. Unlike approaches that rely on linearization approximations, our methodology precisely captures system nonlinearities while incorporating realistic sensor limitations. The proposed framework addresses fundamental challenges in existing multi-rotor controllers by providing formal robustness guarantees against parametric uncertainties and exogenous disturbances, thereby bridging theoretical control design with practical implementation constraints.
@@ II-A Dynamics
We consider the classical multi-rotor UAV dynamics derived from the Newton-Euler equations. Since this work focuses on robust stabilization, the analysis is restricted to the rotational dynamics of the vehicle, while the translational motion is omitted. The nonlinear equations of motion for the rotational dynamics are given by [9]:
$\begin{aligned} \begin{bmatrix}\dot{\phi}\\ \dot{\theta}\\ \dot{\psi}\end{bmatrix}=\begin{bmatrix}\cos{\theta}&0&\sin{\theta}\\ \sin{\theta}&1&-\cos{\theta}\\ -\sin{\theta}\sec{\phi}&0&\cos{\theta}\sec{\phi}\end{bmatrix}\begin{bmatrix}p\\ q\\ r\end{bmatrix}\end{aligned}$, (1)
@@ II-C Robust Control Problem as an LFT Interconnection
where $x$ is the state vector, $y_{m}$ is the measured output, and $w$ is the exogenous input that includes sensor noise and disturbance. $z$ represents the regulated output, which needs to be minimized, which is the attitude and actuator efforts. The vector $\rho$ represents the LPV parameters given in (4). The system matrices are nonlinear functions of $\rho$.
The control objective is to design an output-feedback controller $K$ that regulates $z$ based on the available measurement $y_{m}$, while ensuring robust stability and performance against parameter variations $\rho$ and exogenous disturbances $w$. This is achieved by embedding the system into the generalized interconnection shown in Fig. 1, where weighting functions $W_{i}$ map the physical design specifications into the $\mathcal{H_{\infty}}$ framework.
These weights shape the frequency response capture actuator bandwidth limits, noise spectra, and disturbance rejection requirements that are not directly visible in the raw signals. Moreover, the exogenous inputs are normalized as unit-norm bounded, ensuring that the $\mathcal{H}_{\infty}$ optimization problems remain well-posed, with performance guarantees scaling proportionally with disturbance magnitudes.
@@ III-B Simulation Setup
$w(t)=[d_{\phi}(t)\quad d_{\theta}(t)\quad d_{\psi}(t)\quad n_{p}\quad n_{q}\quad n_{r}]^{\top}$ (9)
where $d_{\phi,\theta,\psi}(t)$ denote disturbance torques induced by wind gusts along the respective axes, and $n_{p,q,r}(t)$ represent sensor noise in gyroscopic measurements.
@@
For baseline comparison, a cascaded SISO PID controller is implemented for each axis, owing to its popularity. The gains are tuned using MATLAB’s autotune feature to ensure best-case PID performance. The $\mathcal{H}_{\infty}$ controller is evaluated against this benchmark.
@@ III-C Results and Discussions
The closed-loop attitude trajectories under PID and $\mathcal{H}_{\infty}$ controllers are presented in Fig. 4-5. Since it is a regulation problem, the controllers reject disturbances and drive the states to zero using only gyroscope measurements. The results demonstrate that both robust controllers achieve significantly superior performance compared to the PID controller, showing improved resilience against disturbances, uncertainties, and measurement noise. In all cases, the errors remain bounded, validating the robustness of the proposed controllers against dynamic and measurement uncertainties.
@@
The Euler angle deviations’ peak and root mean square error (RMSE) values are summarized in Table III. The PID controller exhibits large deviations, with a peak error exceeding $30^{\circ}$, whereas the $\mathcal{H}_{\infty}$ controller reduces the error by a significant order of magnitude.
@@
The control inputs required to achieve the desired states are shown in Fig. 6. The $\mathcal{H}_{\infty}$ controller achieves lower tracking errors without exerting aggressive or high-magnitude actuator efforts. Additionally, the actuator model in the simulation ensures that the controller does not exceed actuator bandwidths. Therefore, the $\mathcal{H}_{\infty}$ controller exhibits overall improved performance relative to the PID controller as it regulates the states effectively with less actuator activity.
Figure 6: Required Control Inputs

### v2

# Robust Attitude Control of Nonlinear UAV Dynamics with LFT Models and $\mathcal{H}_{\infty}$ Performance

@@ I Introduction
A key advantage of robust control frameworks lies in their compatibility with Linear Parameter-Varying (LPV) modeling, which enables exact representation of nonlinear dynamics without local linearization or approximation [12]. The rotational dynamics of multi-rotors contain trigonometric and rational expressions that can be modeled precisely using static nonlinear operators, which are then treated as structured, bounded uncertainty blocks within the Linear Fractional Transformation (LFT) formalism. This representation maintains the full fidelity of the nonlinear dynamics while enabling systematic controller synthesis via the $\mathcal{H}_{\infty}$ technique.
@@ I-A New Contributions
This work studies the nonlinear flight dynamics of a multirotor aerial vehicle within a structured uncertainty framework. The equations of motion considered in this paper correspond to the standard nonlinear rigid-body dynamics commonly used for generic UAV platforms. The formulation is not specific to any particular vehicle configuration and it represents a general multirotor model. The numerical parameter values used in simulations correspond to a representative vehicle for evaluation. Building on this modeling framework, this work advances the application of structured uncertainty frameworks in UAV control by integrating LPV-LFT modeling with robust synthesis techniques. Unlike approaches that rely on linearization approximations, our methodology precisely captures system nonlinearities while incorporating realistic sensor limitations. The proposed framework addresses fundamental challenges in existing multi-rotor controllers by providing formal robustness guarantees against parametric uncertainties and exogenous disturbances, thereby bridging theoretical control design with practical implementation constraints.
@@ II-A Dynamics
We consider the classical multi-rotor UAV dynamics derived from the Newton-Euler equations. Since this work focuses on robust stabilization, the analysis is restricted to the rotational dynamics of the vehicle, while the translational motion is omitted. The nonlinear equations of motion for the rotational dynamics of a multi-rotor UAV are given in [9] with ZYX Euler sequence:
$\begin{aligned} \begin{bmatrix}\dot{\phi}\\ \dot{\theta}\\ \dot{\psi}\end{bmatrix}=\begin{bmatrix}1&\sin{\phi}\tan{\theta}&\cos{\phi}\tan{\theta}\\ 0&\cos{\phi}&-\sin{\phi}\\ 0&\sin{\phi}\sec{\theta}&\cos{\phi}\sec{\theta}\end{bmatrix}\begin{bmatrix}p\\ q\\ r\end{bmatrix}\end{aligned}$, (1)
@@ II-C Robust Control Problem as an LFT Interconnection
where $x$ is the state vector, $y_{m}$ is the measured output, and $w$ is the exogenous input that includes sensor noise and disturbance. $z$ represents the regulated outputs that need to be minimized, which in our case is the attitude error and actuator efforts. The vector $\rho$ represents the LPV parameters given in (4). The system matrices are nonlinear functions of $\rho$.
The control objective is to design an output-feedback controller $K$ that regulates $z$ based on the available measurement $y_{m}$, while ensuring robust stability and performance against parameter variations $\rho$ and exogenous disturbances $w$. This is achieved by embedding the system into the generalized interconnection shown in Fig. 1, where weighting functions $W_{r,d,u,n}$ map the physical design specifications into the $\mathcal{H_{\infty}}$ framework.
These weights shape the frequency response and capture actuator bandwidth limits, noise spectra, and disturbance rejection requirements that are not directly visible in the raw signals. Moreover, the exogenous inputs are normalized as unit-norm bounded, ensuring that the $\mathcal{H}_{\infty}$ optimization problems remain well-posed, with performance guarantees scaling proportionally with disturbance magnitudes.
@@ III-B Simulation Setup
$w(t)=[\tau_{\phi}(t)\quad\tau_{\theta}(t)\quad\tau_{\psi}(t)\quad n_{p}\quad n_{q}\quad n_{r}]^{\top}$ (9)
where $\tau_{\phi,\theta,\psi}(t)$ denote disturbance torques induced by wind gusts along the respective axes, and $n_{p,q,r}(t)$ represent sensor noise in gyroscopic measurements.
@@
For baseline comparison, a cascaded SISO PID controller is implemented for each axis, owing to its popularity. The gains are tuned using MATLAB’s autotune feature to ensure best-case PID performance. This autotuner works by performing a frequency-response estimation experiment to identify the system dynamics, and then calculating PID gains based on desired performance metrics. The $\mathcal{H}_{\infty}$ controller is evaluated against this benchmark.
@@ III-C Results and Discussions
The closed-loop attitude trajectories under PID and $\mathcal{H}_{\infty}$ controllers are presented in Fig. 4-5. Since it is a regulation problem, the controllers reject disturbances and drive the states to zero using only gyroscope measurements. The results demonstrate that the robust controller achieve significantly superior performance compared to the PID controller, showing improved resilience against disturbances, uncertainties, and measurement noise. In all cases, the errors remain bounded, validating the robustness of the proposed controllers against dynamic and measurement uncertainties.
@@
The Euler angle deviations’ peak and root mean square error (RMSE) values are summarized in Table III. The PID controller exhibits large deviations, with a peak error exceeding $30^{\circ}$, whereas the $\mathcal{H}_{\infty}$ controller reduces the error by nearly an order of magnitude.
@@
The control inputs required to achieve the desired states are shown in Fig. 6. The $\mathcal{H}_{\infty}$ controller achieves lower tracking errors without exerting aggressive or high-magnitude actuator efforts. Additionally, the actuator model in the simulation ensures that the controller does not exceed actuator bandwidths. Therefore, the $\mathcal{H}_{\infty}$ controller exhibits overall improved performance relative to the PID controller as it regulates the states effectively with less actuator exertion.
Figure 6: Control Inputs
We also verified that the LPV parameters $\rho(t)$ remain within the assumed bounds under closed-loop operation. Since roll and pitch are constrained to $\phi,\theta\in[\frac{-\pi}{2},\frac{\pi}{2}]$, the trigonometric scheduling terms (e.g., $\sin{(.)},\cos{(.)}$) are naturally bounded in [1,1][-1,1]. The states $p,q,r$ (corresponding to $\operatorname{\rho_{5}-\rho_{7}}$) remain within the prescribed limits ($\pm 1.5\pi$ rad) throughout the simulations, as confirmed by the state trajectories shown in Fig. 5. No violations of the $\rho$ bounds were observed for the $\mathcal{H}_{\infty}$ controller.